OpenStudy (anonymous):

Use l'Hopital's rule to find the limit Limx->infinity: xsin(10/x) Please explain in full detail, I do not know the hopital rule that well.

OpenStudy (anonymous):

you can watch this video for help on hopital rule ( http://www.khanacademy.org/math/calculus/differential-calculus/v/introduction-to-l-hopital-s-rule)

OpenStudy (tkhunny):

Really big hint. \(x\cdot\sin\left(\dfrac{10}{x}\right) = \dfrac{\sin\left(\dfrac{10}{x}\right)}{\dfrac{1}{x}} = 10\cdot\dfrac{\sin\left(\dfrac{10}{x}\right)}{\left(\dfrac{10}{x}\right)}\)

OpenStudy (anonymous):

just make x approach 0 instead, and change the function into sin(x/10)/x = 1/10

OpenStudy (anonymous):

\[xsin(10/x)\] \[\sin \frac{ 10/x }{ x }\] \[10\cos(10(0))\] \[10\]

OpenStudy (anonymous):

is 10 the derivative...??

OpenStudy (anonymous):

which is the derivative??

OpenStudy (anonymous):

\[\frac{ d }{ dx } x \sin(\frac{ 10 }{ x }) = -\frac{ 10 } {x} \cos (\frac{ 10 }{ x })+\sin(\frac{ 10 }{ x })\]

OpenStudy (anonymous):

Thanks !!

OpenStudy (anonymous):

why is it positive 10 shouldn't it be -10

OpenStudy (anonymous):